OpenStudy (horsegirl27):

Please help me out, will medal!! I'm really confused. It's creating a radical equation. I do not understand radical equations at all.

OpenStudy (horsegirl27):

http://prntscr.com/7exg84

OpenStudy (horsegirl27):

Could someone explain radical equations, then help me make 2 for this question? And what does extraneous mean?

OpenStudy (horsegirl27):

@phi @ParthKohli @mathmate

OpenStudy (anonymous):

1) Isolate the radical symbol on one side of the equation 2) Square both sides of the equation to eliminate the radical symbol 3) Solve the equation that comes out after the squaring process 4) Check your answers with the original equation to avoid extraneous values

OpenStudy (horsegirl27):

How do I isolate it?

OpenStudy (phi):

I think I would start by picking a number and squaring it. for example 5 and 5*5= 25 and write \[ \sqrt{25} = 5 \]

OpenStudy (horsegirl27):

Okay.

OpenStudy (phi):

they want \[ a \sqrt{x+b} + c = d\] so far we have \[ \sqrt{25} = 5\] we could write 25 as the sum of two numbers can you pick a pair of numbers that add up to 25 ?

OpenStudy (horsegirl27):

16 and 9?

OpenStudy (phi):

ok. now let x be one of the numbers (which we pretend we don't know) and b be the other number. example: let x be 16 then we can write \[ \sqrt{x+9}= 5 \] (and we know x=16 would make this true) now we want to multiply by "a" \[ a\sqrt{x+9}= a\cdot 5 \] we can pick any number we want for a pick a number for a

OpenStudy (horsegirl27):

3

OpenStudy (phi):

If it is not clear, we are creating a complicated equation that matches what they want \[ a \sqrt{x+b} + c = d \]

OpenStudy (horsegirl27):

I understand how we're replacing numbers for the letters

OpenStudy (phi):

so if we pick a to be 3, then \[ a\sqrt{x+9}= a\cdot 5 \\ 3\sqrt{x+9}= 3\cdot 5 \\ 3\sqrt{x+9}=15\]

OpenStudy (phi):

almost done. they want us to add a number "c" to the left side (and to keep things equal) we do the same to the right side \[ 3\sqrt{x+9} + c=15+c \]

OpenStudy (horsegirl27):

okay, should we use any number for c?

OpenStudy (phi):

yes, we are making it up. It can be anything (but 0 would be cheating)

OpenStudy (horsegirl27):

Okay let's use 7.

OpenStudy (phi):

\[ 3\sqrt{x+9} + c=15+c \\ 3\sqrt{x+9} + 7=15+7 \\ 3\sqrt{x+9} + 7=22 \]

OpenStudy (phi):

I don't know if this has an extraneous solution but I do know x=16 works because that is how we started... which means when you solve it , you better get x=16 as an answer

OpenStudy (horsegirl27):

Do you know what extraneous means?

OpenStudy (phi):

Extraneous means "extra and wrong" to solve this equation we have to undo all the steps we used to make it. to do that, we square the radical , and we get a complicated equation that has more than 1 solution (one of them we do not want)

OpenStudy (phi):

Let's try to solve your equation \[ 3\sqrt{x+9} + 7=22 \] First add -7 to both sides

OpenStudy (horsegirl27):

that will take out +7 and change 22 to 15

OpenStudy (horsegirl27):

\[3\sqrt{x+9}=15 \]

OpenStudy (phi):

now divide both sides by 3

OpenStudy (horsegirl27):

how do I do that? I know 15 will become 5, but what about the left side?

OpenStudy (phi):

I guess you have to know how fractions work. With just numbers: \[ \frac{3 \cdot 2 }{3} = \frac{3}{3} \cdot 2 = 2 \] If the 2 were something else (such as \( \sqrt{x+9} \) it still works the same way

OpenStudy (horsegirl27):

ohhh ok

OpenStudy (phi):

so you get \[ \frac{3 \sqrt{x+9}}{3} = \frac{15}{3} \]

OpenStudy (phi):

which simplifies to ?

OpenStudy (horsegirl27):

Wait, after simplifying, if I divide 3 by 3, will that be changing it to 1?

OpenStudy (phi):

on the left , the 3/3 = 1 on the right the 15/3 = 5

OpenStudy (horsegirl27):

ok, I knew the right side

OpenStudy (phi):

yes, 3/3 =1 and 1 times anything is the anything

OpenStudy (horsegirl27):

So it will be \[1\sqrt{x+3}=5\] or is that wrong?

OpenStudy (phi):

btw, notice \[ \frac{15}{3} = \frac{3 \cdot 5}{3} = \frac{3}{3} \cdot 5 =1\cdot 5= 5 \]

OpenStudy (phi):

You can only divide one thing. if you have \[ \frac{3\cdot 3 \cdot 3}{3} = \frac{3}{3} \cdot 3 \cdot 3 \] notice you don't divide the other 3's. so in our problem the \(\sqrt{x+9} \) is not divided by 3 if that makes sense?

OpenStudy (horsegirl27):

Ohhhh yes it finally makes sense now!

OpenStudy (phi):

another way to keep it straight is to think of dividing as multiplying by the reciprocal in other words \[ \frac{3 \cdot 3 }{3} \text{ means } \frac{1}{3} \cdot 3 \cdot 3 \] (change divide by 3 into multiply by 1/3) it should be more clear you can do \[ \left( \frac{1}{3} \cdot 3 \right) \cdot 3 \]

OpenStudy (horsegirl27):

ok

OpenStudy (phi):

and definitely not \[ \left( \frac{1}{3} \cdot 3 \right) \cdot\left( \frac{1}{3} \cdot 3 \right) \]

OpenStudy (phi):

but back to the problem. what do you have so far?

OpenStudy (horsegirl27):

um I think \[1\sqrt{x+9}=5\]

OpenStudy (phi):

yes, but generally we don't bother to show the multiply by 1

OpenStudy (horsegirl27):

ok

OpenStudy (phi):

in other words we know 1 times \(\sqrt{x+9}\) is \(\sqrt{x+9}\) so we just write \[ \sqrt{x+9}=5 \]

OpenStudy (horsegirl27):

ok, that makes sense

OpenStudy (phi):

we could write (goofy) things like \[ \sqrt{x+9}= 1\cdot 5+ 0 \] but we get into the habit of writing as few symbols as possible

OpenStudy (horsegirl27):

Yeah, we don't need to write all of that

OpenStudy (phi):

to "get rid" of the radical operator, square both sides what do we get?

OpenStudy (horsegirl27):

well, 9 squared is 81, and 5 squared is 25

OpenStudy (phi):

you have \[ \sqrt{x+9}=5 \] "square both sides" means do \[ \left(\sqrt{x+9}\right)^2 =5^2\]

OpenStudy (horsegirl27):

ohhhhhhhhhhhhhhhhhhhhhhhh

OpenStudy (phi):

square means "multiply by itself one time" \[ \left(\sqrt{x+9}\right)\left(\sqrt{x+9}\right) =5 \cdot 5 \]

OpenStudy (horsegirl27):

that makes sense now lol

OpenStudy (phi):

by definition of square root \[ \sqrt{z} \cdot \sqrt{z} = \sqrt{z\cdot z}= z \] it does not matter how "complicated" z is (in our case (x+9) )

OpenStudy (horsegirl27):

ok

OpenStudy (phi):

or, just remember, squaring a square root undoes the square root

OpenStudy (phi):

examples: \[ \left( \sqrt{3x}\right)^2 = 3x \\ \left( \sqrt{x^2+y^2}\right)^2 = x^2 + y^2 \]

OpenStudy (horsegirl27):

that sort of makes sense

OpenStudy (phi):

It is a definition

OpenStudy (horsegirl27):

Okay, reading it a few times, I understand

OpenStudy (horsegirl27):

You square a square root so it's no longer a square root.

OpenStudy (phi):

yes

OpenStudy (horsegirl27):

ok, and then if I was squaring \[\sqrt{x+9}\] it will become just x+9, right?

OpenStudy (phi):

yes and because we have an equation we also square the right side to keep things equal

OpenStudy (horsegirl27):

ok, and what's that going to be? 25, since it wasn't in a square root?

OpenStudy (phi):

yes \[ \left(\sqrt{x+9}\right)^2 =5^2 \\ x+9= 25 \]

OpenStudy (horsegirl27):

Then I do -9, to get x=16

OpenStudy (horsegirl27):

-9 on both sides

OpenStudy (phi):

yes. so that means we have a radical equation \[ 3\sqrt{x+9} + 7=22 \] that does not have an extraneous solution (it has a "true" solution) to make an equation with an extraneous solution, start over with \[ \sqrt{25} = -5 \] notice if we square both sides we get \[ \left( \sqrt{25}\right)^2 = -5 \cdot -5 \\ 25 = 25 \] because -5 , along with +5, is a square root of 25

OpenStudy (horsegirl27):

ok

OpenStudy (horsegirl27):

So it doesn't matter if something is negative or positive to be a square root?

OpenStudy (phi):

a real number has two square roots, and you might see \[ \pm \sqrt{x} \] to remind us but to continue do the same thing as before let 25 be x+9 let a=3, and multiply both sides by 3

OpenStudy (phi):

\[ \sqrt{25} = -5 \\ \sqrt{x+9 } = -5\\3 \sqrt{x+9 } = -5\cdot 3 \]

OpenStudy (phi):

and add 7 to both sides to get \[3 \sqrt{x+9 } +7 = -8 \]

OpenStudy (horsegirl27):

ok

OpenStudy (phi):

can you solve this radical equation ?

OpenStudy (horsegirl27):

I think so, let me work on it

OpenStudy (horsegirl27):

So, first do -7 on both sides and get \[3\sqrt{x+9}=-15\] right? then divide by 3 the same way as before and get \[\sqrt{x+9}=-5\] and then square it to get x+9=25 again. Right?

OpenStudy (phi):

yes, looks good

OpenStudy (horsegirl27):

ok, so it's basically the same.

OpenStudy (phi):

yes, you get x=16 now "put in 16" for x in the original equation. what do you get ?

OpenStudy (horsegirl27):

the original equation being \[3\sqrt{x+9}+7=15?\]

OpenStudy (phi):

this one \[ 3 \sqrt{x+9 } +7 = -8 \]

OpenStudy (horsegirl27):

ok

OpenStudy (phi):

put in 16 for x on the left side, and simplify the left side. what do you get ?

OpenStudy (horsegirl27):

I'm not quite sure how I should simplify it

OpenStudy (phi):

the left side is \[ 3 \sqrt{x+9 } +7 \] replace x with 16: \[ 3 \sqrt{16+9 } +7 = -8 \] order of operations: do the stuff in side the radical first

OpenStudy (phi):

in other words add 16+9

OpenStudy (horsegirl27):

ok, that's 25.

OpenStudy (phi):

yes, you now have \[ 3 \sqrt{25 } +7 \] now do the square root next

OpenStudy (horsegirl27):

5

OpenStudy (phi):

ok, so it's now \[ 3 \cdot 5 +7 \] order of operations: multiply before add

OpenStudy (horsegirl27):

15+7 22

OpenStudy (phi):

yes, and we just found \[ 3 \sqrt{x+9 } +7 = -8 \] with x= 16 becomes \[ 22 = -8 \]